OpenStudy (anonymous):

The square of a number is smaller than 12 times the number by 32. Find the number.

OpenStudy (anonymous):

factor it what did you get?

OpenStudy (dumbcow):

N^2 = 12N -32 N^2 -12N +32 = 0 (N-8)(N-4) = 0 N = 8 or 4

OpenStudy (anonymous):

find x

OpenStudy (anonymous):

(12 - X) * (11 - X) = (132 - 112) = 132 - 23X + X^2 = 20 = X^2 - 23X + 112 X = 16 OR 7 ...16 is not possible .. hence x = 7

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

7